### v2

# A demonstration that classical gravity does not produce entanglement

Abstract
Once again, dispute has arisen over the interpretation of proposed quantum information theory experiments to probe the quantum nature of gravity by testing for gravitationally induced entanglement (GIE) between two spatially separated massive particles ([aziz2025classical] vs. [marletto2025classical, marlettoETC2025classical]; further contributions in [diosi2025no, lin2025newtonian]). The confusion appears to reside in interpreting applications of a Hamiltonian formalism. But classical gravity cannot mediate entanglement on independent grounds. A Newton-Cartan analysis shows that if gravity is classical, a mediator, and entanglement is observed as an outcome of performing a GIE experiment, something other than gravity must have supplied the (virtual) force needed during the experiment to produce the effect.
1 Introduction
In [huggett2023quantum], we discussed at length a recent proposed “GIE” experiment protocol based on quantum information theory to test whether gravity, construed as a mediator between spatially separated massive quantum systems in the laboratory that are otherwise entirely screened off from each other, is itself quantum. Confirmation follows if those otherwise isolated massive quantum systems entangle. The logic of this confirmation depends on a background claim from quantum information theory:
  • Claim
    Interactions with classical systems, particularly mediators governed by classical gravity, do not produce entanglement between quantum systems.
In [aziz2025classical], the authors dispute Claim. They purport to show that classical gravity can produce entanglement by providing a particular Hamiltonian, which they interpret as generating quantum communication and therefore violating an underlying LOCC assumption familiar in quantum information theory (because it violates the CC part: classical communication). Likewise, in the original version (v.1) of [lin2025newtonian], the authors purported to show that classical gravity can produce entanglement by providing a particular Hamiltonian, which they interpret as generating non-local effects and therefore violating the same LOCC assumption (now because it violates the LO part: local operations). In fact, the latter invokes what we explicate in [huggett2023quantum] as a naive “Newtonian model” of GIE: an idealization in which the experiment is described as a bipartite system, with a direct, non-local interaction term, not a tripartite system that includes a gravitational mediator. So such a model is simply silent on any remark, like Claim, which assumes a gravitational mediator. [^5] After posting the initial version of this note to the ArXiv, we learned that shortly prior, a subsequent version of [lin2025newtonian] had reversed the authors’ original pronouncement. In the later version, the authors provide demonstrations in alignment with the standard reasoning that holds Newtonian gravity to be a mediator: “We show that the mesoscopic quantum bodies with their parities of mass quadrupoles interacting via Newtonian gravity can get entangled only if the parity of the gravitational tidal field sourced by the quantum body is also quantized” (v.2 abstract).
[aziz2025classical] (and [lin2025newtonian, v.1]) proceed by interpreting interaction Hamiltonians, and they accept a standard idealization that Newtonian gravity suffices to capture the role of gravity in the experiment protocol, rather than relativistic gravity. Objections to the original authors’ analysis of the Hamiltonian they provide have already appeared [marletto2025classical, marlettoETC2025classical, diosi2025no, di2025simple]. But classical gravity cannot produce entanglement on independent grounds, having nothing to do with a Hamiltonian formalism. In this note, we reiterate our demonstration that in the agreed regime of the experiment, if classical gravity is taken to be the Newtonian limit of classical Einstein gravity, then gravitational mediation simply will not induce entanglement. Claim is on solid grounds.
In what follows, we draw heavily on material that was previously relegated to an appendix B in [huggett2023quantum]. Our impression is that appendices can go unnoticed.
2 A Newton-Cartan Analysis
A central issue appears to be that working in a Hamiltonian formalism encourages taking the Newtonian limit to be a gravitational potential in flat, Galilean spacetime, a view that seriously distorts the nature of gravitational mediation in General Relativity (GR). In many applications this idealization of course does not matter, but if one is specifically concerned with the question of what can be achieved by relativistic gravitational mediation in the Newtonian limit, then it matters a great deal. Fortunately there is an alternative way of taking the limit, which yields the Newton-Cartan formulation—what is arguably more faithful to the nature of general relativity ([malament1986newtonian, christian1997exactly] and references therein). Of course, the more usual flat spacetime formulation of gravity in the Newtonian limit can be obtained from the Newton-Cartan formulation, but Newton-Cartan gravity (NCG) provides a more fruitful and physically appropriate setting to ask and answer foundational questions.
As further evidence for (and explication of) the claim that NCG captures the Newtonian limit of classical gravitational mediation, note that NCG can be quantized into a gauge field theory, whose quanta (“Newton-Cartan gravitons”) are infinite speed longitudinal bosons [christian1997exactly]. These, like the more familiar gravitons of quantized linear GR, form a third subsystem mediating gravity between a pair of massive subsystems. As we pointed out in [huggett2023quantum] this tripartite subsystem picture is essential for sensibly invoking LOCC, so must be respected. The operative sense of ‘local’ is that there is no direct interaction between the masses, but only with the mediator (indeed, our original point was that NCG shows that instantaneous propagation in a classical setting is not sufficient for mediating entanglement). That such a tripartite model arises from quantizing NCG is evidence that one can also safely take classical NCG to capture the Newtonian limit in the most perspicuous way for investigating classical mediation.666[dibiagio2025circuitlocalityrelativisticlocality, di2025simple] argue that LOCC (and other no-go theorems) require a stricter sense of ‘mediator’, which excludes gauge theories, including general relativity, and potentially NCG. We do not claim that our definition of ‘mediator’ is sufficient for LOCC (though it is necessary), only that it is sufficient for a more general notion of mediation, in line with what is expressed by the impressionistic equation 2 in [di2025simple]. Indeed, we take the discussion that closes [dibiagio2025circuitlocalityrelativisticlocality, §F] as tacit acknowledgment that the more general notion is (sometimes) what is physically relevant.
[…]
In [huggett2023quantum, Appendix B], we reproduce the standard entanglement prediction of the GIE experiment protocol working with classical spacetimes. The original purpose was to show a parallel analysis to an existing derivation that uses relativistic geometry, which was provided by [christodoulou2019possibility], to clarify that it is merely the encoding of gravity as spacetime curvature that suffices to generate the effect. One does not need to invoke fundamentality arguments, that deep down we might presume that spacetime satisfies relativistic locality, and that it is the ultimately quantum nature of this latter structure that produces entanglement. In other words: an assumption that gravity is a mediator—in line with the meaning of ‘local’ in ‘LOCC’—and therefore that its mediating is what produces the entanglement effect, makes sense even starting from a manifestly non-local theory like Newtonian gravity, which in other formalisms (namely, a Hamiltonian formalism) includes a direct interaction term.
Whereas in the relativistic setting, the entanglement effect is recognized as a gravitational redshift, in our analysis the entanglement effect is due to a force term over the time of the experiment, which makes up a mismatch between different derivative operators relevant in describing different branches of the wavefunction within the experiment. That is, assuming that basis states in a quantum gravity theory fit for this regime are well approximated as classical spacetimes (in short: that the experiment is in a Newtonian regime), the GIE experiments principally involve two different models in the neighborhood of some curve (with endpoints) $\gamma:[s_{1},s_{2}]\rightarrow M$ modeling the worldline of one of the two massive quantum systems—idealized as a test particle of mass $m$ located at its own center of mass—through the duration of the experiment. These two models, $(M,t_{a},h^{ab},\nabla)$ and $(M,t_{a},h^{ab},\nabla^{\prime})$, differ at most in their curved derivative operators $\nabla$ and $\nabla^{\prime}$.
On the assumption that $\gamma$ is geodesic with respect to $\nabla$, in one of the models but not the other, work is required for the particle to follow that trajectory. In other words, one of the branches of the wavefunction involves (virtual) work for the particle to stick along $\gamma$. As noted in a footnote in that appendix, one may regard this quantity of work as the result of an ordinary force field on $(M,t_{a},h^{ab},\nabla)$, which acts only on particles there that would survey an errant $\nabla^{\prime}$. (If $\nabla^{\prime}=\nabla$ up to a constant multiple, corresponding to a gravitational potential $\phi^{\prime}=\phi$ up to an additive constant, this force field trivializes to the 0 tensor and zero work is put in. In this respect, $\nabla^{\prime}$ is not errant, after all.)
The upshot of this is that the work put in within one branch of the wavefunction and not the other is no more or less mysterious than that which is put into any charged particle in the presence of a fixed, ambient field, which deflects that charged particle off of its geodesic path in accordance with a force law associated with the field-charge pair.
To move from work to total energy over the lifetime of the experiment, one integrates the scalar product of power and proper time over the curve $\gamma$. Thus, if $\xi^{a}$ is the four-velocity of the point mass $m$, here is the expression for the total energy input to the particle over the lifetime of the experiment, with respect to $(M,t_{a},h^{ab},\nabla)$, which is required due to its surveying an errant $\nabla^{\prime}$:
$\text{Total Energy}=m\int_{\gamma}{\xi^{b}(\xi^{m}\xi^{n}t_{m}t_{n})\nabla_{b}(\phi^{\prime}-\phi)(t_{c}\xi^{c})dS}.$ (1)
$t_{c}\xi^{c}$ is the proper time, which in this setting necessarily agrees with the absolute time elapsed between the two endpoints of $\gamma$, the total lifetime of the experiment $T$. Given the static nature of the experiment protocol, the expression further simplifies as $m(\phi^{\prime}-\phi)*T$, where $(\phi^{\prime}-\phi)$ is understood as a function solely of spatial coordinates at a point, corresponding to the difference between the gravitational potentials associated with $\nabla^{\prime}$ and $\nabla$. Finally, as discussed in that appendix, this quantity can be regarded as equivalent to the value of a classical action, which upon quantization yields a phase factor in agreement with existing analyses of the experiment.
What is significant for present purposes, however, is that in the case where the underlying theory is not a quantum theory whose basis states are approximated as distinct models, and instead there is just one single model (whatever it is), everything in the preceding analysis goes through, but with $\nabla^{\prime}=\nabla$. It is immediate that the phase factor trivializes (disregarding that computing the phase factor is arguably no longer physically motivated if the wavefunction branches are identical), and we can clearly say why: the relevant force field that showed up earlier in the analysis is the 0 tensor. To produce entanglement requires otherwise.
Note that while we have argued that NCG is the appropriate Newtonian limit for the analysis of classical gravitational mediation (hence our demonstration that it induces no entanglement), there are outstanding questions regarding the applicability of our quantum Newton-Cartan derivation of the standard phase difference. First, one would like to recover this picture from the quantized NCG developed in [christian1997exactly], in analogy to the relativistic case [chen2023quantum]. Second, there is the much harder question of whether quantized NCG as formulated in [christian1997exactly] can be obtained as the Newtonian limit of some more plausible theory of quantum gravity than quantized linear GR, given that classical NCG is the limit of full GR.
3 Commentary
Our simple results establish that classical gravity, formulated as NCG, a manifestly non-local theory, cannot as a mediator induce entanglement. We stress that all parties agree that, within the regime of the experiment protocol, this particular theory is the appropriate treatment of classical Einstein gravity. The use of the Newton-Cartan formalism just makes clear what is at stake in an antecedent commitment to think of classical gravity as a mediator (despite its non-local character as that mediator).
What this means is that if entanglement is observed as an outcome of performing a GIE experiment, and if gravity is classical, something else is responsible for the entanglement.
Acknowledgements
We thank Andrea di Biagio for helpful explications on the content of his note [di2025simple].